INFLUENCE OF CHEMICALLY ACTIVE ADDITIVES ON EMISSION SPECTRA OF THE FLAMES OF HYDROGEN AND DEUTERIUM OXIDATION IN NEAR IR REGION

\underline{N. M. RUBTSOV}, G. I. TSVETKOV and V. I. CHERNYSH, Institute for Structural Macrokinetics and Material Science of Russian Academy of Sciences, 142432, Russian Federation, Noginski region, Chernogolovka.

Band system in the region 852 − 880 nm and bands 670 and 846 nm over pressures 0.1 − 1 atm in the flames of mixtures of hydrogen with oxygen and air has been detected. The system is assigned to overtone transitions of H_{2}O_{2} (3ν_{1} + σ) correspondingly arising from the reaction of HO_{2}+HO_{2}−H_{2}O_{2}^{v}+O_{2}+33.5 kcal/mole. The bands 670 and 846 nm are assigned to overtone transitions of HO_{2} radical (4ν_{1} + σ) and (3ν_{1} + σ) correspondingly arising from the reaction of H+O_{2}(+M)−HO_{2}(+M) + 44 kcal/mole. By this means the extent of energy of these elementary processes is located in vibrational degrees of freedom of products. It has been detected that HO_{2} radicals occur originally in overtone-excited states in combustion both of hydrogen and deuterium in the presence of small additives of propylene as an inhibitor.